# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 5. A homeomorphism of the pseudoarc that is not conjugate to an automorphism of the pre-pseudoarc
  • and construct epimorphisms $h_{i}:J_{n_{i}+1}\to J_{n_{i}}$ in such a way that for any choice of $\tilde{\sigma}_{i}\in$ Aut($\mathbb{P}$}) with $\pi_{n_{i}}\circ\tilde{\sigma}_{i}=h_{i}\circ\pi_{n_{i}+1}$ we have that $\tilde{\sigma}_{i}$ converges uniformly (in ${\mathcal{C}}(P,P)$), moreover the limit is an element of ${\rm Homeo}(P)$,
@@
Suppose that $f:J\to L$ is an epimorphism between the finite linear graphs $J$, $L$, and let $a\in J$ be a node, $M\subseteq J$ be a subinterval with $a$ being an endpoint of $M$. Then we define the type of the pair $a,L$ with respect to $f$, in symbols, ${\rm tp}^{a,M}(f)$ to be the $L$-type
@@ 5. $\blacksquare_{5}$
    $\pi_{i,i+1}[K_{i+1}]=K_{i}$, and the $\pi$’s agree with the $h_{i}$$h_{i}^{\bullet}$$g_{i}$$f_{i}$ on $K_{k}$ whenever it is appropriate (and defined), i.e.
@@ 6. $\blacksquare_{6}$
    for each $a,b\in J_{i+1}$, if $aR^{3}b$, then $\pi_{i,i+1}(a)R\pi_{i,i+1}(b)$ (e.g. if for $c\in J_{i}$, each connected component in $\pi_{i,i+1}^{-1}(c)$ is an interval of length at least $3$),
@@ 4. $\blacksquare_{9}(d)$
      for each maximal connected component (subinterval, of which there are $3$) $C$ of $J_{4k+2}\setminus(K_{4k+2}\cup L_{4k+2})$ we have
@@ 3. $\blacksquare_{10}(c)$
      letting $C_{1},C_{2},C_{3}$ denote the connected components in $J_{4k+2}\setminus(K_{4k+2}\cup L_{4k+2})$ we have that letting
@@
      $C_{1}^{\prime},C_{2}^{\prime},C_{3}^{\prime}$ are all connected (intervals), and
@@ 6. $\blacksquare_{10}(f)$
      for each $a\in L_{4k+3}$ we have $|f_{k}(a)-f^{\prime}_{k}(a)|\leq 2$, and if $a\in L_{4k+3}$ is such that $\pi_{4k+2,4k+3}(a)\in L_{4k+2}$ is neither an endpoint nor is related to an endpoint of $L_{4k+2}$, then
      $|f_{k}(a)-f^{\prime}_{k}(a)|=2\ (=|\pi_{4k+2,4k+3}(a)-f^{\prime}_{k}(a)|)$
@@
      (5.7) $(\pi_{4k+1,4k+2}\circ f_{k})R(g_{k}\circ\pi_{4k+2,4k+3}),$
@@ 3. $(3)$
    there exists $(x_{i})_{i}$, $(y_{i})_{i}\in K$ such that $\pi_{i}((x_{i})_{i})=x_{i}$, $\pi_{i}((y_{i})_{i})=y_{i}$ are the two endpoints of $K_{j}$,
@@ 4. $(4)$
Since $K_{i}$ is an interval, any clopen decomposition $C_{0}\cup C_{1}$ of $K$ is of the form $C_{0}=\pi_{i}^{-1}(K^{0}_{i})$, $C_{1}=\pi_{i}^{-1}(K^{1}_{i})$, for a partition $K^{0}_{i}\cup K^{1}_{i}$ of $K_{i}$. Now there must exist $(z^{0}_{i})_{i},(z^{1}_{i})_{i}\in K$ with $z^{0}_{i}\in K^{0}_{i}$, $z^{1}_{i}\in K^{1}_{i}$ and $(z^{0}_{i})_{i}R(z^{1}_{i})_{i}$. But then necessarily ${\rm pr}((z^{0}_{i})_{i})={\rm pr}((z^{1}_{i})_{i})$, so $C$ is a continuum, indeed.
For $(3)$ note that $\{x_{i},y_{i}\}$ must be the set of the two endpoints of $K_{i}$. Now clause $\blacksquare_{5}$ clearly implies $(3)$.
Pick such an $\sigma_{k}$ for each $k$. First we are going to check that $(\sigma_{k})_{k}$ is convergent in $\mathcal{C}(P,P)$ (so $\lim_{k}\sigma_{k}$ is a surjective continuous function), and that $\lim_{k}\sigma_{k}$ is injective. In light of the second part of Theorem 2.4 it suffices to verify that for all finite connected linear graph $A$ and $\varphi\in{\rm Epi}({\mathbb{P}},A)$ we have that for all large enough $n,n^{\prime}$ we have $(\varphi\circ\sigma_{n})R(\varphi\circ\sigma_{n^{\prime}})$. Using the fact that every epimorphism factors through a $\pi_{j}$, it is enough to check that
(5.8) $(\pi_{4k-1}\circ\sigma_{k})R(\pi_{4k^{\prime}-1}\circ\sigma_{k^{\prime}})\ \textrm{ if }k\leq k^{\prime}.$
@@
(Claim 5.8) First we note that since $\pi_{4k+3}\circ\sigma_{k+1}=h_{k+1}\circ\pi_{4k+4}$, clause $\blacksquare_{7}$$\blacksquare_{7}(b)$ implies
(5.9) $\begin{array}[]{l}\pi_{4k+2}\circ\sigma_{k+1}=\pi_{4k+2,4k+3}\circ h_{k+1}\circ\pi_{4k+4}=f_{k}\circ\pi_{4k+3,4k+4}\circ\pi_{4k+4}\\ =f_{k}\circ\pi_{4k+3}.\end{array}$
As (5.7) says
$(\pi_{4k+1,4k+2}\circ f_{k})R(g_{k}\circ\pi_{4k+2,4k+3}),$
we can apply $\pi_{4k+1,4k+2}$ to (5.9)
(5.10) $\pi_{4k+1}\circ\sigma_{k+1}=\pi_{4k,4k+2}\circ f_{k}\circ\pi_{4k+3}\ R\ (g_{k}\circ\pi_{4k+2}).$
On the other hand, applying $\pi_{4k+1,4k+3}$ to $\psi R(\pi_{4k+3}\circ\sigma_{k+1})$ we get
$(\pi_{4k+1,4k+3}\circ\psi)R(\pi_{4k+1}\circ\sigma_{k+1}),$
so by (5.10),
$(g_{k}\circ\pi_{4k+2})R^{2}(\pi_{4+1,4k+3}\circ\psi).$
@@
but notice that this LHS is just $\pi_{4k-1}\circ\sigma_{k}$, and we are done. ∎
@@
(Claim 5.9) Pick $(z_{i})_{i}$, $(z^{\prime}_{i})_{i}$ in $\mathbb{P}$} with ${\rm pr}((z_{i})_{i})\neq{\rm pr}((z^{\prime}_{i})_{i})$. This means that $\neg(z_{i}Rz_{i}^{\prime})$ for large enough $i$. With (5.8) it is enough to show that for large enough $k$ we have $\neg(\pi_{4k-1}\circ\sigma_{k}((z_{i})_{i})R^{3}\pi_{4k-1}\circ\sigma_{k}((z^{\prime}_{i})_{i})$, i.e (by the definition of $\sigma_{k}$), we need that
@@
(5.12) $(\pi_{4k+2,4k+3}\circ h_{k+1}(z_{4k+4}))R(\pi_{4k+2,4k+3}\circ h_{k+1}(z^{\prime}_{4k+4})$
by $\blacksquare_{6}$. Applying $\blacksquare_{7}$$\blacksquare_{7}(b)$ to the LHS,
(5.13) $(\pi_{4k+2,4k+3}\circ h_{k+1}(z_{4k+4}))=(f_{k}\circ\pi_{4k+3,4k+4})(z_{4k+4})=f_{k}(z_{4k+3}),$
and similarly for $z^{\prime}_{4k+4}$, i.e.
(5.14) $(\pi_{4k+2,4k+3}\circ h_{k+1}(z^{\prime}_{4k+4}))=(f_{k}\circ\pi_{4k+3,4k+4})(z^{\prime}_{4k+4})=f_{k}(z^{\prime}_{4k+3}).$
Combining (5.12)-(5.14) we get $(f_{k}(z_{4k+3})R(f_{k}(z^{\prime}_{4k+3}),$ so clearly
(5.15) $(\pi_{4k+1,4k+2}\circ f_{k}(z_{4k+3})R(\pi_{4k+1,4k+2}\circ f_{k}(z^{\prime}_{4k+3}).$
Applying (5.7),
(5.16) $(\pi_{4k+1,4k+2}\circ f_{k}(z_{4k+3}))R(h_{4k+1}\circ\pi_{4k+2,4k+3})(z_{4k+3})=g_{k}(z_{4k+2}),$
and similarly for $z^{\prime}$, i.e.
(5.17) $(\pi_{4k+1,4k+2}\circ f_{k}(z^{\prime}_{4k+3}))R(h_{4k+1}\circ\pi_{4k+2,4k+3})(z^{\prime}_{4k+3})=g_{k}(z^{\prime}_{4k+2}),$
and so (5.15)-(5.17) give us $g_{k}(z_{4k+2})R^{3}g_{k}(z^{\prime}_{4k+2})$. Then by $\blacksquare_{8}$$\blacksquare_{8}(c)$
@@
By ($\blacksquare_{9}$$\blacksquare_{9}(e)$), $h^{\bullet}_{k}\circ g_{k}(z_{4k+2})=z_{4k}$, and $h^{\bullet}_{k}\circ g_{k}(z^{\prime}_{4k+2})=z^{\prime}_{4k}$, so $z_{4k}Rz_{4k}^{\prime}$, which contradicts (5.11). ∎
So $\sigma\in{\rm Homeo}(P)$, indeed, and it remains to check that $C$, $x$, $y$ witness that $\sigma$ has the properties from Lemma 5.2. First we recall Fact A.2, i.e. that sets of the form
@@
Now we can fix the pairwise disjoint regular open sets $U_{1},U_{2},U_{3}$ of $P$ as in Lemma 5.2 with $x\in U_{2}$, $y\in U_{3}$, $\bigcup_{i=1}^{4}{\rm cl}(U_{i})=P$ and $C\cap{\rm cl}(U_{1})=\emptyset$. Let $W_{i}={\rm pr}^{-1}(U_{i})$ for $i=1,2,3$. By the observation above there exists $i_{0}$ such that
@@ 2. $\boxdot_{5}$
    $\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-1})\cup\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-2})\cup\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-3})\subseteq W_{3}$
which we can assume by possibly flipping the order and numbering (and in fact this is not even strict, but we won’t need more than these 33-33 elements). Now we recall $\blacksquare_{10}$$\blacksquare_{10}(f)$, i.e. for $a\in L_{4k+3}$ we have $|f_{k}(a)-\pi_{4k+2,4k+3}(a)|=2$ unless $\pi_{4k+2,4k+3}(a)\in\{l_{0},l_{1},l_{|L_{4k+2}|-1},l_{|L_{4k+2}|-2}\}$. This can only happen if either

### v2

@@ 1. $\sigma{\restriction}C={\rm id}_{C}$;
  • and construct an increasing sequence $(n_{i})_{i}$ and epimorphisms $h_{i}:J_{n_{i}+1}\to J_{n_{i}}$ in such a way that for any choice of $\tilde{\sigma}_{i}\in$ Aut($\mathbb{P}$}) with $\pi_{n_{i}}\circ\tilde{\sigma}_{i}=h_{i}\circ\pi_{n_{i}+1}$ we have that $\tilde{\sigma}_{i}$ converges uniformly (in ${\mathcal{C}}(P,P)$), moreover the limit is an element of ${\rm Homeo}(P)$,
@@
Suppose that $f:J\to L$ is an epimorphism between the finite linear graphs $J$, $L$, and let $a\in J$ be a node, $M\subseteq J$ be a subinterval with $a$ being an endpoint of $M$. Then we define the type of the pair $a,M$ with respect to $f$, in symbols, ${\rm tp}^{a,M}(f)$ to be the $f[M]$-type
@@
Basically the functions $h_{i}^{\prime}$, $f_{i}$, $g_{i}$ are liftings of each other, the only purpose to give them different names is an attempt to ease the notational awkwardness later, e.g. in the proof of Lemma 5.6. More concretely, $h_{i}:J_{4i}\to J_{4i-1}$ is an epimorphism and $h^{\bullet}_{i}:J_{4i+1}\to J_{4i}$ is an epimorphism with $h_{i}\circ h^{\bullet}_{i}=\pi_{4i-1,4i+1}$, where the existence of $h^{\bullet}_{i}$’s will ensure that the limit is (left-)invertible, so injective. Then $g_{i}:J_{4k+2,4k+1}$ will be a lifting of $h_{i}$, which will be ensured by the condition $h^{\bullet}_{i}\circ g_{i}=\pi_{4i,4i+2}$. In the recursive construction $K_{4i+1}$ will not only have $K_{4i+2}$ as a $g_{i}$- and $\pi_{4i+1,4i+2}$-preimage (on which $g_{i}$ and $\pi_{4i+1,4i+2}$ coincide), but also a copy of it $L_{4i+2}$. In the next step of the recursion this is followed by the construction of $J_{4i+3}$, and $f_{i},\pi_{4i+2,4i+3}:J_{4i+3}\to J_{4i+2}$, where we can also guarantee that $f_{i},\pi_{4i+2,4i+3}$ map $L_{4i+3}$ to $L_{4i+2}$, as well as $K_{4i+2}$ to $K_{4i+2}$. But instead of $f_{i}$ being a lifting of $g_{i}$, it will be only an almost lifting, because we will arrange so that $f_{i}{\restriction}L_{4i+3}$ is roughly speaking a shift of $\pi_{4i+3,4i+2}{\restriction}L_{4i+3}$ by one. (This is the crucial step to guarantee that regular open partitions are never $\sigma$-invariant.) Finally, $h_{i+1}:J_{4i}\to J_{4i-1}$ will be a lifting of $f_{i}$ guaranteed by the condition $f_{i}\circ\pi_{4i+3,4i+4}=\pi_{4i+2,4i+3}\circ h_{i+1}$. Each $h_{i}$ will be responsible for an amalgamation task to ensure genericity of the sequence. We will have further technical conditions, e.g. the conditions on types, which will be necessary to carry out our main tasks.
@@ 5. $\blacksquare_{5}$
    $\pi_{i,i+1}[K_{i+1}]=K_{i}$, and the $\pi$’s agree with the $h_{i}$$h_{i}^{\bullet}$$g_{i}$$f_{i}$ on $K_{k}$ whenever it is appropriate (and defined), that is,
@@ 6. $\blacksquare_{6}$
    for each $a,b\in J_{i+1}$, if $aR^{3}b$, then $\pi_{i,i+1}(a)R\pi_{i,i+1}(b)$ (e.g. this is ensured if for $c\in J_{i}$, each connected component in $\pi_{i,i+1}^{-1}(c)$ is an interval of length at least $3$),
@@ 4. $\blacksquare_{9}(d)$
      for each of the three maximal connected component (interval) $C$ in $J_{4k+2}\setminus(K_{4k+2}\cup L_{4k+2})$ we have
@@ 3. $\blacksquare_{10}(c)$
      letting $C_{1},C_{2},C_{3}$ denote the connected components in $J_{4k+2}\setminus(K_{4k+2}\cup L_{4k+2})$ if we let
@@
      then we have that $C_{1}^{\prime},C_{2}^{\prime},C_{3}^{\prime}$ are all connected (intervals), and
@@ 6. $\blacksquare_{10}(f)$
      for each $a\in L_{4k+3}$ we have ${\rm dist}(f_{k}(a),f^{\prime}_{k}(a))\leq 2$, and if $a\in L_{4k+3}$ is such that $\pi_{4k+2,4k+3}(a)\in L_{4k+2}$ is neither an endpoint nor is related to an endpoint of $L_{4k+2}$, then
      ${\rm dist}(f_{k}(a),f^{\prime}_{k}(a))=2\ (={\rm dist}(\pi_{4k+2,4k+3}(a),f^{\prime}_{k}(a))$
@@
      (5.20) $(\pi_{4k+1,4k+2}\circ f_{k})R(g_{k}\circ\pi_{4k+2,4k+3})$
      (here ${\rm dist}(a,b)\leq n$ means $aR^{n}b$),
@@ 3. $(3)$
    there exists $(x_{i})_{i}$, $(y_{i})_{i}\in K$ such that $\pi_{j}((x_{i})_{i})=x_{j}$, $\pi_{j}((y_{i})_{i})=y_{j}$ are the two endpoints of $K_{j}$,
@@ 4. $(4)$
Similarly, by a routine argument it follows from the fact that the $K_{i}$’s are interval that $C$ is a continuum, but for the sake of completeness we elaborate. Since $K_{i}$ is an interval, any clopen decomposition $C_{0}\cup C_{1}$ of $K$ is of the form $C_{0}=\pi_{i}^{-1}(K^{0}_{i})$, $C_{1}=\pi_{i}^{-1}(K^{1}_{i})$, for a partition $K^{0}_{i}\cup K^{1}_{i}$ of $K_{i}$. Now there must exist $(z^{0}_{i})_{i},(z^{1}_{i})_{i}\in K$ with $z^{0}_{i}\in K^{0}_{i}$, $z^{1}_{i}\in K^{1}_{i}$ and $(z^{0}_{i})_{i}R(z^{1}_{i})_{i}$. But then necessarily ${\rm pr}((z^{0}_{i})_{i})={\rm pr}((z^{1}_{i})_{i})$, so $C$ is a continuum, indeed.
For $(3)$ note that $\{x_{i},y_{i}\}$ can only be the set of the two endpoints of $K_{i}$. Now clause $\blacksquare_{5}$ clearly implies $(3)$.
Pick such an $\sigma_{k}$ for each $k$. First we are going to check that $(\sigma_{k})_{k}$ is convergent in $\mathcal{C}(P,P)$ (so $\lim_{k}\sigma_{k}$ is a surjective continuous function), and that $\lim_{k}\sigma_{k}$ is injective. In light of the second part of Theorem 2.4 it suffices to verify that for all finite connected linear graph $A$ and $\varphi\in{\rm Epi}({\mathbb{P}},A)$ we have that for all large enough $n,n^{\prime}$ the relation $(\varphi\circ\sigma_{n})R(\varphi\circ\sigma_{n^{\prime}})$ holds. Using the fact that every epimorphism factors through a $\pi_{j}$, it is enough to check that
(5.21) $(\pi_{4k-1}\circ\sigma_{k})R(\pi_{4k-1}\circ\sigma_{k^{\prime}})\ \textrm{ if }k\leq k^{\prime}.$
@@
(Claim 5.7) Using that $\pi_{4k+3}\circ\sigma_{k+1}=h_{k+1}\circ\pi_{4k+4}$ and $\pi_{4k-1}\circ\sigma_{k}=h_{k}\circ\pi_{4k}$ we need to argue that if $\psi$ is $R$-related to $h_{k+1}\circ\pi_{4k+4}$, then the composition $(\pi_{4k-1,4k+3}\circ\psi)$ is $R$-related to $h_{k}\circ\pi_{4k}$.
We are going to argue that (roughly speaking) as $h_{k+1}$ is a lifting of $f_{k}$, and $f_{k}$ is an almost lifting of $g_{k}$ (within distance one), so $(g_{k}\circ\pi_{4k+2})$ and $(\pi_{4k+1,4k+3}\circ\psi)$ are within distance two. At this point we will compose these two with $\pi_{4k,4k+1}$, use the contractive nature of these mappings together with the fact that $g_{k}$ is a lifting of $h_{k}$.
More concretely we first argue that
(5.22) $(\pi_{4k+1,4k+3}\circ h_{k+1})R(g_{k}\circ\pi_{4k+2,4k+4}),$
which will then immediately imply
(5.23) $(g_{k}\circ\pi_{4k+2})R(\pi_{4k+1,4k+3}\circ h_{k+1}\circ\pi_{4k+4})R(\pi_{4k+1,4k+3}\circ\psi).$
To see (5.22), we note that clause $\blacksquare_{7}$$\blacksquare_{7}(b)$ implies
(5.24) $\pi_{4k+1,4k+3}\circ h_{k+1}=\pi_{4k+1,4k+2}\circ f_{k}\circ\pi_{4k+3,4k+4}$
and (5.20) implies
$(\pi_{4k+1,4k+2}\circ f_{k}\circ\pi_{4k+3,4k+4})R(g_{k}\circ\pi_{4k+2,4k+4}),$
so (5.22) holds, indeed.
@@
and we are done. ∎
@@
(Claim 5.8) Pick $(z_{i})_{i}$, $(z^{\prime}_{i})_{i}$ in $\mathbb{P}$} with ${\rm pr}((z_{i})_{i})\neq{\rm pr}((z^{\prime}_{i})_{i})$. This means that $\neg(z_{i}Rz_{i}^{\prime})$ for large enough $i$. With (5.21) in order to argue that $\sigma({\rm pr}((z_{i})_{i}))\neq\sigma({\rm pr}((z^{\prime}_{i})_{i}))$ it is enough to show that for large enough $k$ we have
$\neg(\pi_{4k-1}\circ\sigma_{k}((z_{i})_{i}))\ R^{3}\ (\pi_{4k-1}\circ\sigma_{k}((z^{\prime}_{i})_{i})),$
, that is, (by the definition of $\sigma_{k}$), we need that
@@
(5.26) $(\pi_{4k+1,4k+3}\circ h_{k+1}(z_{4k+4}))R(\pi_{4k+1,4k+3}\circ h_{k+1}(z^{\prime}_{4k+4}))$
by $\blacksquare_{6}$.
We are going to argue that approximating the $\pi_{4k+1,4k+3}$-image of $h_{k+1}(z_{4k+4})$ and $h_{k+1}(z_{4k+4})$ by $g_{k}(z_{4k+2})$ and $g_{k}(z^{\prime}_{4k+2})$, the latter two are $R^{3}$-related, to which we can apply $h^{\bullet}_{k}$, and use $h^{\bullet}_{k}\circ g_{k}=\pi_{4k,4k+2}$ to get a contradiction with (5.25).
To formalize this idea we use (5.22) to get
(5.27) $(\pi_{4k+1,4k+3}\circ h_{k+1}(z_{4k+4}))R(g_{k}\circ\pi_{4k+2,4k+4})(z_{4k+4})=g_{k}(z_{4k+2}),$
and similarly for $z^{\prime}$, that is,
(5.28) $(\pi_{4k+1,4k+3}\circ h_{k+1}(z^{\prime}_{4k+4}))R(g_{k}\circ\pi_{4k+2,4k+4})(z^{\prime}_{4k+4})=g_{k}(z^{\prime}_{4k+2}).$
So combining(5.26)-(5.28) we obtain
$g_{k}(z_{4k+2})R^{3}g_{k}(z^{\prime}_{4k+2}).$
Then by $\blacksquare_{8}$$\blacksquare_{8}(c)$
@@
By ($\blacksquare_{9}$$\blacksquare_{9}(e)$), $h^{\bullet}_{k}\circ g_{k}(z_{4k+2})=z_{4k}$, and $h^{\bullet}_{k}\circ g_{k}(z^{\prime}_{4k+2})=z^{\prime}_{4k}$, so clearly $z_{4k}Rz_{4k}^{\prime}$, which contradicts (5.25). ∎
So $\sigma\in{\rm Homeo}(P)$, indeed, and it remains to check that $C$, $x$, $y$ witness that $\sigma$ has the properties from Lemma 5.2. First we recall Proposition A.3, that is, that sets of the form
@@
Now we can fix the pairwise disjoint regular open sets $U_{1},U_{2},U_{3}$ of $P$ as in Lemma 5.2 with $x\in U_{2}$, $y\in U_{3}$, $\bigcup_{i=1}^{3}{\rm cl}(U_{i})=P$ and $C\cap{\rm cl}(U_{1})=\emptyset$. Let $W_{i}={\rm pr}^{-1}(U_{i})$ for $i=1,2,3$. By the observation above there exists $i_{0}$ such that
@@ 2. $\boxdot_{5}$
    $(\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-1})\cup\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-2})\cup\pi^{-1}_{4k+2}(l_{|L_{4k+2}|-3}))\cap(W_{1}\cup W_{2})=\emptyset$
which we can assume by possibly flipping the order and numbering (and in fact this is not even strict, but we won’t need more than these 33-33 elements). We introduce the informal notation
$l_{n}\dotplus m=l_{n+m}$
if $0\leq n,n+m<|L_{4k+2}|$. Now we recall $\blacksquare_{10}$$\blacksquare_{10}(f)$, that is, for $a\in L_{4k+3}$ we have ${\rm dist}(f_{k}(a),\pi_{4k+2,4k+3}(a))=2$ unless $\pi_{4k+2,4k+3}(a)\in\{l_{0},l_{1},l_{|L_{4k+2}|-1},l_{|L_{4k+2}|-2}\}$. This can only happen if either